# Graph-based Approaches and Functionalities in Retrieval-Augmented Generation: A Comprehensive Survey

### v1

@@ 5.2.2. Methods Using Convolutional-based Algorithms
REANO (ref:reano, ) is a framework designed to leverage structural information in knowledge graphs to enhance the retrieval of triple features by employing an L-hop graph convolution. Before the retrieval step, REANO uses the TAGME model (ferragina2010tagme, ) to extract entities and then builds relationships among these entities through both intra-context and inter-context manners. Subsequently, an $L$-hop GNN is applied to update entity representations, allowing REANO to select the top-$k$ triple candidates based on their similarity to the given question. By incorporating the topological structure over multiple hops within the knowledge graph, REANO gathers critical information needed to answer questions, effectively alleviating the reasoning burden of the answer predictor and improving the overall efficiency of the QA process.
SURGE (ref:SURGE, ) is a framework designed to extract context-relevant knowledge from knowledge graphs while enforcing consistency across facts to ensure that the generated responses faithfully reflect the retrieved knowledge corresponding to the question. To achieve this, SURGE leverages an existing edge message passing framework (jo2021edge, ), which transforms the edges of the original graph into nodes within a dual hypergraph (scheinerman2011fractional, ). This transformation enables the use of existing node-level GNNs to represent the relationships within the original graph. Subsequently, SURGE samples negative subgraphs and employs contrastive learning, encouraging the model to generate responses that are consistent with the positive subgraph, thereby improving the fidelity of the generated content. By enforcing consistency through contrastive learning, SURGE ensures that its responses remain faithful to the retrieved knowledge, enhancing the reliability of the generated answers.
Explore-then-Determine (EtD) (ref:etd, ) utilizes a graph convolutional method to enhance graph retrieval by employing adaptive propagation and message passing. In the Explore phase, the framework initializes with a topic entity and uses a lightweight GNN, integrated with semantic representations from LLMs, to expand the candidate set by propagating to relevant neighbors. Attention weights are computed to rank and prune irrelevant edges, focusing on the top-$K$ edges at each layer to refine candidate sets. In the Determine phase, the filtered graph is processed by a frozen LLM with a knowledge-enhanced multiple-choice prompt, merging explicit graph-based knowledge with the LLM’s implicit understanding to determine the final answer. This approach efficiently filters irrelevant KG information while enabling precise multi-hop reasoning.
@@ 5.2.3. Methods Using Attention-Based Algorithms
HSGE (ref:hsge, ) studies the complex interactions between the entities of the knowledge base by reasoning in the history semantic graph, which is built by employing a pre-trained language model BERT (ref:bert, ) on conversation contexts. A fundamental method of HSGE to retrieve the structure information in the history semantic graph is to update the node embeddings with the attention-based module such as TransformerConv (TransformerConv, ). In order to further adapt to the change triggered by the new-coming conversations, HSGE encodes the position of each historical interaction and utilizes the attention mechanism again to aggregate the most relevant information for each mentioned entity in the query. Compared with concatenating all previous conversation turns, this paradigm is more computationally efficient, while still retaining key context information needed for effective question answering.
HamQA (ref:hamqa, ) intends to identify the importance of different relationships between the entities by emphasizing the information most relevant for addressing the question. Specifically, HamQA formulates the weights of each edge in the KG utilizing a learnable function combined with the entity representations. Moreover, considering the geometrically hierarchical features between different entities, HamQA further employs Hyperbolic distances (balazevic2019multi, ) to measure the importance of different neighbors. These two strategies enable HamQA to effectively calculate the attention scores when incorporating the adjacent information and serve as the constraints to guide graph propagation toward more significant messages.
@@
Topology-Aware Prompting Preserves structure, enables multi-hop reasoning Complex formatting, hard to understand for LLMs LLaGA (ref:llaga, ), GNN-RAG (ref:gnn-rag, ), G-Retriever (ref:g-retriever, ), GRAPH-COT (ref:graphcot, ), ROG (ref:rog, ), RRA (ref:rra, ), MVP-Tuning (ref:mvp-tuning, ), Keqing (ref:keqing, ), GRAG (ref:grag, ), ETD (ref:etd, ), DALK (ref:dalk, ), MindMap (ref:mindmap, ), Lark (ref:lark, ), REALM (zhu2024realm, ), HybridRAG (ref:hybridrag, ), EWEK-QA (ref:ewek-qa, ), TOG2 (ref:tog2, ), SURGE (ref:SURGE, ), GraphGPT (ref:GraphGPT, ), NLGraph (ref:nlgraph, ), GraphEval2000 (ref:GraphEval2000, ), MuseGraph (ref:musegraph, )
Text Prompting Simple implementation, compatible with LLM input format Loses structured information, limited reasoning ability Talk (ref:talk, ), GPT4Graph (ref:GPT4Graph, ), UniOQA (ref:unioqa, ), FABULA (ref:fabula, ), ODA (ref:oda, ), KnowGPT (ref:knowgpt, ), Kaping (ref:kaping, ), MedGraphRAG (ref:medgraphrag, ), Golden-Retriever (ref:golden, ), KALMV (ref:kalmv, ), Fact (ref:fact, ), DepsRAG (ref:depsrag, ), KELP (ref:kelp, ), KGQA (ref:kgqa, ), Graph-LLM (ref:Graph-LLM, ), GLBench (ref:GLBench, ), Walklm (ref:walklm, )
@@ 5.4.1. A Unified View
Topology-aware prompting captures the topological essence of a knowledge graph by explicitly encoding nodes, edges, and their relationships into structured prompts that models capable of directly processing structured inputs can readily interpret. Instead of providing simple textual descriptions, these prompts explicitly represent relationships using structured formats such as triple statements (e.g., ”(Albert Einstein, born in, Ulm)”) or relational paths (e.g., ”Paris → capital of → France → located in → Europe”). A key motivation behind graph-structure prompting is to allow the model to reason more effectively about multi-hop and complex relational patterns (ref:tog2, ). For instance, rather than treating the graph as a mere collection of facts, the model can examine path-based relationships to draw inferences. This is especially valuable in scenarios where the desired answer depends on understanding connections among multiple entities or interpreting intricate graph substructures. Moreover, this approach can boost explainability by making the model’s reasoning process more transparent. When each relationship or path is clearly defined, users can trace how the system derived its conclusions, thereby increasing trust and clarity in the model’s output. In summary, graph-structure prompting offers a powerful mechanism for harnessing a graph’s rich relational data, allowing RAG-based systems to deliver more reliable answers and deeper insights.
@@ 5.4.2. Methods Applying the Topology-Aware Prompting
Keqing (ref:keqing, ) utilizes graph-structured prompting to enhance KGQA tasks by guiding LLMs through complex multi-hop questions. It achieves this by employing two distinct prompting strategies to help LLMs understand and reason about the logical relationships in retrieved triplets from a KG. The first strategy explicitly explains the triplet structure in the prompt, detailing the format as (subject, relation, object), and ensuring that responses align strictly with the entities within the triplets. The second strategy converts triplets into plain text by serializing their components, making them more digestible for fine-tuned models like LLaMA(llama, ). These graph-structured prompts act as a chain-of-thought (COT) mentor for LLMs, enabling them to decompose complicated queries into simpler sub-questions, retrieve relevant entities through logical chains, and generate accurate and interpretable responses. By leveraging the structured nature of KGs and tailored prompting techniques, Keqing demonstrates improved reliability and interpretability in KGQA tasks, highlighting its potential as a scalable solution for knowledge-intensive reasoning.
GRAG (ref:grag, ) introduces a dual approach to graph-structure prompting for integrating textual graphs into LLMs. In the Graph View, textual graphs are encoded as soft prompts using a Graph Neural Network (GNN) to preserve topological information. Retrieved subgraph embeddings are aligned with LLM token embeddings via an MLP, ensuring seamless integration of graph structure. In the Text View, textual graphs are transformed into hierarchical text descriptions as hard prompts. This involves splitting retrieved ego-graphs into a Breadth-First Search (BFS) tree and residual edge set, then organizing them into structured templates to retain narrative and topological context. The query, hierarchical text description, and graph embeddings are concatenated as input to the LLM, enabling context-aware generation that leverages both structural and semantic aspects of the graph.
MindMap (ref:mindmap, ) creates prompts for LLMs by converting knowledge graph subgraphs into natural language descriptions. First, it mines evidence subgraphs using path-based and neighbor-based exploration techniques to gather relevant information. Then, it performs an evidence graph aggregation step, where each subgraph is formatted as an entity chain, such as $\mathcal{G}^{\prime}=\{(e_{h},r,e_{t})\}$, which is translated into a natural sentence by a predefined template. These descriptions are merged to form a reasoning graph, which is included as part of the LLM’s prompt. This approach ensures that the LLM understands both the structural and semantic aspects of the graph, enabling robust and context-aware reasoning.
KGP (ref:kgp, ) formulates the process of generating knowledge graph prompts for multi-document question answering scenarios into two steps. First, a general knowledge graph is constructed based on the passage similarity from multiple documents. During the process, the merits and limitations of different construction methods are examined, to build the mapping between the most appropriate methods and scenarios. Second, a language model agent is employed to traverse the knowledge graph and retrieve relevant context. The agent traverses over the KG through the new evidence given by the LLM iteratively according to the retrieved passages. The final structured passage queue from the KG services acts as the prompt for multi-document question answering.
The Explore-then-Determine (EtD) (ref:etd, ) framework employs a multiple-choice prompting method to enhance reasoning over knowledge graphs. This prompt is designed to evaluate and compare candidate answers derived from the graph, incorporating key elements to guide the decision-making process. It includes a task description, the query, and the top-N ranked candidate entities with their associated probabilities, providing a confidence measure for each option. Additionally, evidence chains connecting the topic entity to each candidate are integrated into the prompt, showcasing the compositional relationships between entities. These components enable the model to leverage both structured knowledge and contextual evidence, ensuring accurate and interpretable responses.
@@
Both RoG (ref:rog, ) and ToG (ref:tog, ) generate reasoning paths from KGs to enhance RAG prompting by structuring retrieval and reasoning. RoG focuses on generating relation-grounded paths, ensuring faithful and structured reasoning by retrieving valid reasoning paths from KGs. This approach enhances accuracy and consistency while providing interpretable results that improve trust and understanding. Similarly, ToG identifies relevant entities from a query and explores the KG by searching for meaningful triples. During reasoning, the LLM evaluates the retrieved triples and selects the most valuable ones to construct a reasoning path. This method offers explicit and editable reasoning paths, improving explainability and allowing users to trace and correct model outputs when necessary. Both approaches strengthen RAG by integrating structured retrieval with reasoning, ensuring more coherent, interpretable, and contextually relevant responses.
@@ 5.5.2. Methods Applying the Text Prompting
MedGraphRAG (ref:medgraphrag, ) is a specialized framework that adapts existing LLMs to the medical domain, emphasizing safety and dependability in processing sensitive medical information. To effectively represent the complex relationships among entities in medical knowledge, MedGraphRAG utilizes the LLMs to extract entities from medical documents and constructs relationships between them using source and definition chunks. For efficient information retrieval, MedGraphRAG identifies the top-k relevant entities based on the concatenated embedding similarity with user queries. The system then iteratively refines its answers by feeding them back into the LLMs until it reaches a predefined threshold epoch. This approach allows the LLMs to build a thorough understanding by engaging with all pertinent data in the graph, while maintaining efficiency by summarizing less pertinent information. Ultimately, MedGraphRAG delivers accurate responses in a resource-conscious manner, ensuring effective knowledge retrieval within a high-stakes domain.
KCQA (ref:kgqa, ) builds a framework that efficiently retrieves and prompts knowledge from a KG to answer questions. Initially, KCQA estimates the distribution of relations in the KG using the Rigel model (oliya2021end, ) and subsequently predicts relevant triples within two hops. To focus on the most relevant information, it retains only the top-K triples of the estimated relations, achieved through a Hadamard (element-wise) product to extract a weighted vector representation of the triples. These selected triples are then converted into natural language for prompting. In the prompting process, KCQA also considers inverse triples and utilizes a unified template to compose the prompt, which is then fed as the input into a language model generating an answer.
@@ 5.6. Comparison between Different Prompting Methods
Topology-Aware prompting and text prompting offer fundamentally different ways of integrating graph-based knowledge into RAG systems. Topology-Aware prompting retains the relational dependencies within the graph, enabling precise multi-hop reasoning and enhanced interpretability. However, it relies on models capable of processing structured inputs effectively. In contrast, text prompting re-formats graph information into natural language, improving accessibility but potentially compromising structural clarity. A promising avenue for advancing these techniques is adaptive reasoning calibration (xie2024calibrating, ), where the system dynamically determines the level of structural detail required based on the query’s complexity. Instead of rigidly selecting between structured or textual prompts, models could simplify, expand, or reformat graph-derived knowledge depending on the reasoning depth needed. For instance, straightforward fact retrieval may be optimally handled with natural language, while intricate multi-hop inferences could benefit from structured representations enriched with contextual cues. Additionally, reinforcement learning-driven prompt optimization (huang2024enhancing, ) could enable models to refine their approach over time, learning which formatting strategies yield the most reliable and interpretable results across diverse tasks.
@@ 6.1. A Unified View
Figure 6. Pipeline formulation with graph structure
@@
Sequential Pipeline Simple and efficient execution Fixed order, no refinement or error correction GNN-RAG (ref:gnn-rag, ), GrapeQA (ref:grapeqa, ), SR (ref:sr, ), RRA (ref:rra, ), KG-GPT (ref:kg-gpt, ), QA-GNN (ref:qa-gnn, ), GLBK (ref:noname5, ), ATLANTIC (ref:atlantic, ), UniOQA (ref:unioqa, ), GRAG (ref:grag, ), HyKGE,  (ref:hykge, ), KG-Rank (ref:kg-rank, ), KnowledGPT (ref:knowledgpt, ), ETD (ref:etd, ), HippoRAG (ref:hipporag, ), DALK (ref:dalk, ), KnowGPT (ref:knowgpt, ), Engine (ref:engine, ), GraphBridge (ref:graphbridge, ), ChatKBQA (ref:chatkbqa, ), NuTrea (ref:nutrea, ), etc
@@
Loop Pipeline Iterative refinement, error correction Higher computational cost, low efficiency GraphRAG (ref:graphrag, ), GRAPH-COT (ref:graphcot, ), KnowledgeNavigator (ref:kn, ), KD-COT (ref:kd-cot, ), KG-Agent (ref:kg-agent, ), GNN-Ret (ref:gnn-ret, ), Knowledge Solver (ref:ksl, ), KGP (ref:kgp, ), FABULA (ref:fabula, ), ODA (ref:oda, ), QR-LLM (ref:rewrite1, ), MufassirQAS (alan2024rag, ), ArcaneQA (ref:arcaneqa, ), MedGraphRAG (ref:medgraphrag, ), KALMV (ref:kalmv, )
@@
Tree Pipeline Parallel exploration, multi-path reasoning Higher computational cost EWEK-QA (ref:ewek-qa, ), HamQA (ref:hamqa, ), SURGE (ref:SURGE, ), TOG2 (ref:tog2, )

### v2

@@ 5.2.2. Methods Using Convolutional-based Algorithms
REANO (Fang et al., 2024) is a framework designed to leverage structural information in knowledge graphs to enhance the retrieval of triple features by employing an L-hop graph convolution. Before the retrieval step, REANO uses the TAGME model (Ferragina and Scaiella, 2010) to extract entities and then builds relationships among these entities through both intra-context and inter-context manners. Subsequently, an $L$-hop GNN is applied to update entity representations, allowing REANO to select the top-$k$ triple candidates based on their similarity to the given question. By incorporating the topological structure over multiple hops within the knowledge graph, REANO gathers critical information needed to answer questions, effectively alleviating the reasoning burden of the answer predictor and improving the overall efficiency of the QA process.
SURGE (Kang et al., 2022) is a framework designed to extract context-relevant knowledge from knowledge graphs while enforcing consistency across facts to ensure that the generated responses faithfully reflect the retrieved knowledge corresponding to the question. To achieve this, SURGE leverages an existing edge message passing framework (Jo et al., 2021), which transforms the edges of the original graph into nodes within a dual hypergraph (Scheinerman and Ullman, 2011). This transformation enables the use of existing node-level GNNs to represent the relationships within the original graph. Subsequently, SURGE samples negative subgraphs and employs contrastive learning, encouraging the model to generate responses that are consistent with the positive subgraph, thereby improving the fidelity of the generated content. By enforcing consistency through contrastive learning, SURGE ensures that its responses remain faithful to the retrieved knowledge, enhancing the reliability of the generated answers.
@@ 5.2.3. Methods Using Attention-Based Algorithms
HSGE (Sun et al., 2023) studies the complex interactions between the entities of the knowledge base by reasoning in the history semantic graph, which is built by employing a pre-trained language model BERT (Devlin et al., 2019) on conversation contexts. A fundamental method of HSGE to retrieve the structure information in the history semantic graph is to update the node embeddings with the attention-based module such as TransformerConv (Shi et al., 2020). In order to further adapt to the change triggered by the new-coming conversations, HSGE encodes the position of each historical interaction and utilizes the attention mechanism again to aggregate the most relevant information for each mentioned entity in the query. Compared with concatenating all previous conversation turns, this paradigm is more computationally efficient, while still retaining key context information needed for effective question answering.
HamQA (Dong et al., 2023b) intends to identify the importance of different relationships between the entities by emphasizing the information most relevant for addressing the question. Specifically, HamQA formulates the weights of each edge in the KG utilizing a learnable function combined with the entity representations. Moreover, considering the geometrically hierarchical features between different entities, HamQA further employs Hyperbolic distances (Balazevic et al., 2019) to measure the importance of different neighbors. These two strategies enable HamQA to effectively calculate the attention scores when incorporating the adjacent information and serve as the constraints to guide graph propagation toward more significant messages.
@@
Topology-Aware Prompting Preserves structure, enables multi-hop reasoning Complex formatting, hard to understand for LLMs LLaGA (Chen et al., 2024b), GNN-RAG (Mavromatis and Karypis, 2024), G-Retriever (He et al., 2024), GRAPH-COT (Jin et al., 2024b), ROG (Luo et al., 2024b), RRA (Wu et al., 2023b), MVP-Tuning (Huang et al., 2023b), Keqing (Wang et al., 2023a), GRAG (Hu et al., 2024), ETD (Liu et al., 2024a), DALK (Li et al., 2024a), MindMap (Wen et al., 2024), Lark (Choudhary and Reddy, 2024), REALM (Zhu et al., 2024b), HybridRAG (Sarmah et al., 2024), EWEK-QA (Dehghan et al., 2024), TOG2 (Ma et al., 2024), SURGE (Kang et al., 2022), GraphGPT (Tang et al., 2024), NLGraph (Wang et al., 2024a), HyKGE (Jiang et al., 2024c), HamQA (Dong et al., 2023b), GraphEval2000 (Wu et al., 2024c), MuseGraph (Tan et al., 2024a), GOT (Besta et al., 2024a)
Text Prompting Simple implementation, compatible with LLM input format Loses structured information, limited reasoning ability Talk (Fatemi et al., 2023), GPT4Graph (Guo et al., 2023), UniOQA (Li et al., 2024e), FABULA (Ranade and Joshi, 2023), ODA (Sun et al., 2024b), KnowGPT (Zhang et al., 2024c), Kaping (Baek et al., 2023b), MedGraphRAG (Wu et al., 2024b), Golden-Retriever (An et al., 2024), KALMV (Baek et al., 2023c), Fact (Opsahl, 2024), DepsRAG (Alhanahnah et al., 2024), KELP (Liu et al., 2024b), KGQA (Sen et al., 2023), Graph-LLM (Chen et al., 2024c), GLBench (Li et al., 2024d), Walklm (Tan et al., 2024b)
@@ 5.4.1. A Unified View
Topology-aware prompting captures the topological essence of a knowledge graph by explicitly encoding nodes, edges, and their relationships into structured prompts that models capable of directly processing structured inputs can readily interpret. These featured prompting methods can be divided into following categories: (1) path prompting that encodes relation chains or shortest paths to expose composition (GRAPH-COT (Jin et al., 2024b), ROG (Luo et al., 2024b), TOG2 (Ma et al., 2024), MindMap (Wen et al., 2024)); (2) subgraph prompting that linearizes compact subgraphs for information summarization (G-Retriever (He et al., 2024), DALK (Li et al., 2024a), GRAG (Hu et al., 2024), LLaGA (Chen et al., 2024b)); (3) rule-aware prompting that introduces relation types and declarative constraints to guide reasoning (HyKGE (Jiang et al., 2024c), HamQA (Dong et al., 2023b)).
Instead of providing simple textual descriptions, these prompts explicitly represent relationships using structured formats such as triple statements (e.g., ”(Albert Einstein, born in, Ulm)”) or relational paths (e.g., ”Paris → capital of → France → located in → Europe”). A key motivation behind graph-structure prompting is to allow the model to reason more effectively about multi-hop and complex relational patterns (Ma et al., 2024). For instance, rather than treating the graph as a mere collection of facts, the model can examine path-based relationships to draw inferences. This is especially valuable in scenarios where the desired answer depends on understanding connections among multiple entities or interpreting intricate graph substructures. Moreover, this approach can boost explainability by making the model’s reasoning process more transparent. When each relationship or path is clearly defined, users can trace how the system derived its conclusions, thereby increasing trust and clarity in the model’s output. In summary, graph-structure prompting offers a powerful mechanism for harnessing a graph’s rich relational data, allowing RAG-based systems to deliver more reliable answers and deeper insights.
@@ 5.4.2. Methods Applying the Topology-Aware Prompting
Keqing (Wang et al., 2023a) utilizes graph-structured prompting to enhance KGQA tasks by guiding LLMs through complex multi-hop questions. It achieves this by employing two distinct prompting strategies to help LLMs understand and reason about the logical relationships in retrieved triplets from a KG. The first strategy explicitly explains the triplet structure in the prompt, detailing the format as (subject, relation, object), and ensuring that responses align strictly with the entities within the triplets. The second strategy converts triplets into plain text by serializing their components, making them more digestible for fine-tuned models like LLaMA 2 (Touvron et al., 2023). These graph-structured prompts act as a chain-of-thought (COT) mentor for LLMs, enabling them to decompose complicated queries into simpler sub-questions, retrieve relevant entities through logical chains, and generate accurate and interpretable responses. By leveraging the structured nature of KGs and tailored prompting techniques, Keqing demonstrates improved reliability and interpretability in KGQA tasks, highlighting its potential as a scalable solution for knowledge-intensive reasoning.
@@
Both RoG (Luo et al., 2024b) and ToG (Sun et al., 2024a) generate reasoning paths from KGs to enhance RAG prompting by structuring retrieval and reasoning. RoG focuses on generating relation-grounded paths, ensuring faithful and structured reasoning by retrieving valid reasoning paths from KGs. This approach enhances accuracy and consistency while providing interpretable results that improve trust and understanding. Similarly, ToG identifies relevant entities from a query and explores the KG by searching for meaningful triples. During reasoning, the LLM evaluates the retrieved triples and selects the most valuable ones to construct a reasoning path. This method offers explicit and editable reasoning paths, improving explainability and allowing users to trace and correct model outputs when necessary. Both approaches strengthen RAG by integrating structured retrieval with reasoning, ensuring more coherent, interpretable, and contextually relevant responses.
@@ 5.5.2. Methods Applying the Text Prompting
MedGraphRAG (Wu et al., 2024b) follows a general, reusable pattern of graph-based RAG: graph building, graph retrieval, prompting, and refinement. In the build step, an LLM reads medical documents, extracts entities, links relations, and keeps short evidence or definition snippets so that each node and edge is grounded in source text. At query time, the system embeds the user question and graph entities, selects the top-k most similar entities, and expands a small working subgraph around them for focused context. The model then generates a draft answer from this subgraph and iteratively improves it by feeding the draft back into the LLM for a few rounds, stopping at a preset threshold and summarizing low-value regions to remain efficient. This simple loop—evidence-linked graph construction, embedding-based top-k retrieval with local expansion, and short refinement—captures a general template for reliable, source-aware graph-grounded QA in high-stakes domains.
KCQA (Sen et al., 2023) builds a framework that efficiently retrieves and prompts knowledge from a KG to answer questions. Initially, KCQA estimates the distribution of relations in the KG using the Rigel model (Oliya et al., 2021) and subsequently predicts relevant triples within two hops. To focus on the most relevant information, it retains only the top-K triples of the estimated relations, achieved through a Hadamard (element-wise) product to extract a weighted vector representation of the triples. These selected triples are then converted into natural language for prompting. In the prompting process, KCQA also considers inverse triples and utilizes a unified template to compose the prompt, which is then fed as the input into a language model generating an answer.
@@ 5.6. Comparison between Different Prompting Methods
The strengths and limitations of topology-Aware prompting and text prompting map naturally to task demands. For multi-hop question answering (Tang and Yang, 2024), personalized recommendation (Deldjoo et al., 2024), temporal reasoning (Zhu et al., 2025), and program synthesis over graphs (Gu and Su, 2022), topology-aware prompting is preferable because it preserves edge types and path structure, enforces constraints, and yields auditable chains. However, as shown in Table 3, it brings heavy formatting overhead, larger token budgets, schema canonicalization burdens, and can be harder for general LLMs to parse, making it brittle under noisy graphs (Hadi et al., 2023). By contrast, for abstractive summarization over documents attached to the graph (Edge et al., 2024), for broad retrieval where lexical recall dominates (Choi et al., 2023), and for user-facing explanations that prioritize fluency (Wang et al., 2023c), text prompting is the better choice. It is simple to implement, aligns with standard LLM inputs, remains robust under schema drift, and achieves lower latency, but it sacrifices explicit path identity, weakens constraint satisfaction, and limits compositional reasoning depth (Mao et al., 2024).
@@ 6.1. A Unified View
Figure 6. Pipeline formulation with graph structure
@@
Sequential Pipeline Simple and efficient execution Fixed order, no refinement or error correction GNN-RAG (Mavromatis and Karypis, 2024), GrapeQA (Taunk et al., 2023), SR (Zhang et al., 2022), RRA (Wu et al., 2023b), KG-GPT (Kim et al., 2023a), QA-GNN (Yasunaga et al., 2021), GLBK (Delile et al., 2024), ATLANTIC (Munikoti et al., 2023), UniOQA (Li et al., 2024e), GRAG (Hu et al., 2024), HyKGE,  (Jiang et al., 2024c), KG-Rank (Yang et al., 2024a), KnowledGPT (Wang et al., 2023c), ETD (Liu et al., 2024a), HippoRAG (Gutiérrez et al., 2024), DALK (Li et al., 2024a), KnowGPT (Zhang et al., 2024c), Engine (Zhu et al., 2024c), GraphBridge (Wang et al., 2024c), ChatKBQA (Luo et al., 2024a), NuTrea (Choi et al., 2023), etc
@@
Loop Pipeline Iterative refinement, error correction Higher computational cost, low efficiency GraphRAG (Edge et al., 2024), GRAPH-COT (Jin et al., 2024b), KnowledgeNavigator (Guo et al., 2024), KD-COT (Wang et al., 2023b), KG-Agent (Jiang et al., 2024a), GNN-Ret (Li et al., 2024f), Knowledge Solver (Feng et al., 2023), KGP (Wang et al., 2024d), FABULA (Ranade and Joshi, 2023), ODA (Sun et al., 2024b), QR-LLM (Ma et al., 2023), MufassirQAS (Alan et al., 2024), ArcaneQA (Gu and Su, 2022), MedGraphRAG (Wu et al., 2024b), KALMV (Baek et al., 2023c)
@@
Tree Pipeline Parallel exploration, multi-path reasoning Higher computational cost EWEK-QA (Dehghan et al., 2024), HamQA (Dong et al., 2023b), SURGE (Kang et al., 2022), TOG2 (Ma et al., 2024), GOT (Besta et al., 2024a), Beyond-CoT (Yao et al., 2023)
6.2. Controllers of Pipeline Implementation
The controller is the decision-making mechanism that formalizes our pipeline topologies at inference time, which chooses when to retrieve, whether to iterate or branch, when to verify and revise, and when to stop. Different controller families emphasize sequential, iterative loop, and branching tree pipelines, each with distinct tradeoffs among accuracy, latency, and computational cost. In what follows, we present four widely adopted controller families in RAG systems and show how representative methods instantiate each family within our pipeline design.
Reinforcement-Learning Controllers. A reinforcement-learning (RL) controller is a learned policy that, given the current reasoning state, selects the next pipeline action (retrieve, expand, verify, revise, or stop) to maximize a reward that trades off answer quality and computational cost. For example, RAG-RL (Huang et al., 2025) post-trains the generator with RL so it learns to select and cite useful contexts from larger retrieved sets, which tightens the coupling between retrieval and generation and effectively learns the loop and tree branching rules for exploration and pruning. Multi-module RL has also been used to coordinate several RAG components jointly (query rewriter, retriever, selector, generator), as in MMOA-RAG (Chen et al., 2025), which frames the pipeline as cooperative agents optimized toward a unified reward and thus implements a learned loop controller and branch selection over candidate contexts.
System-2 Slow-Thinking Controllers. A system-2 controller is an inference-time search policy that allocates more computation than a single forward pass, which naturally realizes loop and tree pipelines without updating model weights. For instance, Self-Ask (Press et al., 2022) decomposes a question into sub-questions, interleaves retrieval during the reasoning process, and thus implements a loop that alternates “ask–retrieve–answer” over a constrained subgraph. FLARE (Jiang et al., 2023c) generalizes this idea to long-form generation by forecasting upcoming content, triggering retrieval only when predicted segments are low confidence, and regenerating with fresh evidence.
Verifier-Guided Controllers. A verifier-guided controller attaches an external evaluator that scores partial solutions or retrieved sets and feeds decisions back to the pipeline, thereby governing revision and selection. Self-RAG (Asai et al., 2024) trains a model to retrieve, generate, and critique using reflection tokens. During inference, it adaptively decides whether to retrieve again, accept or revise a segment, and produce citations, which instantiates a loop controller with verify–revise cycles over the current subgraph. CRAG (Yan et al., 2024) adds a lightweight retrieval evaluator that estimates the quality of retrieved documents and triggers corrective actions, such as re-retrieval or web fallback, effectively steering the loop and pruning poor branches in a tree of candidate evidence.
Budget-/Cost-Aware Controllers. A budget-aware controller explicitly optimizes an accuracy–latency objective by selecting per-query computation and retrieval budgets, thereby choosing whether a sequence suffices or whether a loop or tree is warranted.Adaptive-RAG (Jeong et al., 2024) predicts query complexity and selects among no-retrieval, single-step retrieval, or iterative retrieval strategies, implementing a controller that upgrades a sequence pipeline into a loop only when complexity demands. Conversational gating methods such as RAGate (Wang et al., 2024b) estimate, at each turn, whether external augmentation is needed and hence decide to skip or invoke retrieval, providing a cost-aware switch between sequence and loop behavior under compute or latency constraints.